he Oklahoma Pipeline Company projects the following pattern of inflows from an investment. The inflows are spread over time to reflect delayed benefits. Each year is independent of the others. Year 1 Year 5 Year 10 Cash Inflow Probability Cash In

The expected value for all three years is $70.

Compute the standard deviation for each of the three years. (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

Final answer

Standard deviation
Year - 1 13.42
Year - 5 23.24
Year - 10 44.72

Explanation

Year - 1

Let X represents cash flow variables of Year - 1

Dx = Deviation from expected value = X - 70 ( note : expected value is given as 70 )

Dx2 = Dx * Dx

P = Probability

X Dx Dx2 P P * Dx2
55 -15 225 0.4 90
70 0 0 0.2 0
85 15 225 0.4 90
Sum of P * Dx2 = 180
Standard deviation 13.42

Standard deviation = Square Root of ( 180 ) = 13.42 ........... for Year - 1

Year - 5

X Dx Dx2 P P * Dx2
40 -30 900 0.3 270
70 0 0 0.4 0
100 30 900 0.3 270
Sum of P * Dx2 = 540
Standard deviation 23.24

Year - 10

X Dx Dx2 P P * Dx2
20 -50 2500 0.4 1000
70 0 0 0.2 0
120 50 2500 0.4 1000
Sum of P * Dx2 = 2000
Standard deviation 44.72